The sum of three consecutive integers is 72. What is the middle integer ?

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

z = y + 1
y = x + 1
x + y + z = 72
---
put the system of linear equations into standard form
---
y - z = -1
x - y = -1
x + y + z = 72
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 23
y = 24
z = 25
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by Alan3354(69443)

72/3 = 24, the middle integer.
================
n-1 + n + n+1 = 72